PIP-II report for May 4, 2018


  • A kick-off meeting for the design of the Site Preparation work was held on May 3, 2018.
  • Held first PIP-II ESH focused IPT meeting this week with focus on international code equivalencies and contributions from in-kind partners.
  • Defined top-level structure for PIP-II CAD model as the format for incoming lattice data.
  • A wire scanner was installed in the PIP2IT MEBT and a new toroid was installed in the LEBT region.


ESH&Q (Teri Dykhuis & John Anderson):

  • The IDNR and USFWS did not think that a survey for the Rusty Patched Bumble Bee was necessary because it’s known that the species is present. They requested specific information for both the PIP-II and LBNF sites. Our consultant, Arcadis is drafting letters to both agencies that will cover all species and both projects. These letters will be sent from FSO to the agencies.
  • Held first PIP-II ESH focused IPT meeting this week. Focus was on international code equivalencies and contributions from in-kind partners.


Project Management (Marc Kaducak):

  • Scheduled a risk workshop for July 9-10, 2018. The workshop will be facilitated by Keith Molenaar from Colorado State University, who will also facilitate the LBNF/DUNE risk workshop the same week.
  • Eileen Berman from LBNF/DUNE and the Scientific Computing Division introduced the PIP-II team to the LBNF/DUNE document management systems. PIP-II plans to leverage the architecture that LBNF/DUNE has built to enhance commonality across the two projects.
  • The PIP-II Procurement Manager job posting is now open here.


Technical Integration (Arkadiy Klebaner):

  • Advanced CD2 Design Readiness Plan and Design Deliverables Document for CD2;
  • Requirements Verification and Interface Control Documents work proceeding;
  • Assembling Preliminary Design Report table of content;
  • Met with LBNF Systems Engineering team to discuss requirements management, received tutorial on DOORS NG (commercial software for requirements and interfaces management) implementation;
  • Evaluated LBNF/DUNE documentation management strategy as possible framework for PIP-II documentation management;
  • Received tutorial on Teamcenter requirements management tool;
  • Interface development is advancing, with project wide L2 ICDs in initial development and L3 ICDs at 35% completion;
  • Working on Requirements Matrix spreadsheet and a generic safety requirements statement for inclusion into all FRS’s;
  • Started the discussion on the interplay between the L3 FRS documents and the L2 System Configuration Document;


International Coordination (Steve Holmes & Lia Merminga):

  • Confirmed with Department of Commerce/BIS that the existing Fermilab Export License covers sharing of software with BARC.
  • A table of assumed deliverables from international partners was developed. This will be used for building our resource loaded schedule.
  • RFPI Technical Requirements Specification Rev 7 released, and ready for approval by DAE.


SRF & Cryo Systems (Allan Rowe & Genfa Wu):


  • All RF ports outside of the HWR CM changed from SMA to Type-N (female) per Fermilab’s request.
  • Teamcenter workflow procedure for the HWR Interface Specification Document (ISD) is finalized.
  • Assigned L3-managers received the HWR Interface Specification Document (ISD) for review and approval.
  • Preparation for high power integrated (HWR&Coupler) test at ANL is ongoing.


  • SSR1 Cavities
    • Results of cavity S107 tested in STC with LPC were analyzed and cavity meets the requirement.
    • Installation of cavity S111 in the STC. Cooldown to 2K and testing in progress.
    • Cavity S113 received HPC at MP9. Next 120C baking and STC-HPC testing.
    • Flange maintenance of cavity S106. Next cavity to be processed (light BCP+HPR) at ANL.
    • Validated the alignment model of cavities with the use of CMM.
    • Updated tooling list and parts for string assembly with action items (i.e., electropolishing, location, etc.)
    • Updating step-by-step procedure for string assembly.
  • SSR1 Cryomodule
    • Working on completing tooling for WS2-4 (string assembly through final cryomodule) at Lab 2.
    • Drawings of the lifting tooling for WS2 have been checked and released. Next: procurement.
  • Lab2
    • Preparing facility for final cleanroom qualification.


  • The four (2 conventional design + 2 new “shielded” design) coupler cold ends have been received at Fermilab.
  • First 650 MHz coupler meeting held. Current status reviewed, potential action items have been identified. This meeting is proposed to occur on a weekly basis.


  • HB650 1st CM P6 schedule being reviewed and changes being recommended.

Cryogenics Plant

  • Finished draft of ICD
  • Began reviewing Cryoplant building TRS and compiling list of building/utility requirements from Cryoplant perspective.
  • Began compiling list of documentation requirements for Preliminary Design Review of Cryoplant Integration.
  • In contact with ESS cryo folks and scheduled an informal visit to learn various aspects of the ESS cryogenic plant and system.

Cryogenics Distribution System

  • Finalized travel plans to ESS to learn various aspects of the ESS cryogenic plant and systems. The ESS cryogenic system is very similar to the proposed PIP-II cryogenic system.
  • Began compiling list of documentation requirements for Preliminary Design Review of Cryoplant Integration.
  • Completing ICD documents.


Accelerator Systems (Paul Derwent):

High Power RF

  • Argonne is working on installing the AC power for the Tomco amplifier. FNAL Personnel plan to go there for a visit next week.
  • Amplifiers at CMTF have been operating well (RFQ and buncher cavities for PIP2IT).
  • Stakeholders in 650 MHz STC RF distribution have informally agreed on a routing path and installation process for the coax cable run through the HTS-2 enclosure.
  • 2-D drawings of 650 MHz coupler test stand RF distribution are complete.


  • During the PIP2IT 2-day shutdown, the MEBT wire scanner was installed in the beamline. Alignment scheduled for next week.


  • CD-2 work in progress with drafts of the review plan and interface control document complete.
  • PIP2IT RFQ vacuum work continues. During a 2-day shutdown, the pumping port side plate at Module3 right side was removed, inspected, and modified. It has been re-installed and vacuum pumpdown is underway.


Linac Installation & Commissioning (Fernanda G. Garcia):


  • With most of the team participating in IPAC’18, the activities in PIP2IT were mainly focused on installation.
    • RFQ vacuum work
      • This is the primary work for the shutdown.
      • Removal of one port blank off on the right side to assess suspicious o’ring sealing problems.
      • Plate was machined, welded and bench leak checked prior re-installation on the RFQ.
    • LEBT toroid swap
      • A new toroid was installed in the LEBT region to improve signal/noise ratio.
    • LEBT chopper connector shield was installed.


  • Wire scanner installation
    •  Wire scanner was installed, and a rough alignment of the fork was performed.
    • Checks on the limit switch were performed and adjusted accordingly. Also, the full out position to MPS connection was successfully tested.

Test Infrastructure:

  • Preparatory work continues for the PIP2IT cryogenic transfer line.
    • The north run is fully assembled and ready to be installed. Piping is staged at the CMTS building.
    • The east run is being assembled at MW9.


  • Defined top-level structure for PIP2 CAD model as the format for incoming lattice data.


Conventional Facilities (Steve Dixon):

  • A/E Acquisition Plan Corrections Underway – The technical corrections have been made and Procurement will address the contractual comments with a goal of returning the acquisition plan to DOE.
  • Cryo Plant Building Requirements Requisition In Process – The requisition for this work has been circulated for signatures and in Procurement. A purchase order is expected to be issued in the coming week with a start of the work in mid-June 2018.
  • Site Prep Design Kick-off Meeting – A kick-off meeting for the design of the Site Preparation work was held on May 3, 2018.